# The Milky Way as Seen by Classical Cepheids I: Distances Based on Mid-infrared Photometry

### v1

@@ Abstract
Classical Cepheids are the archetype of the standard candle, thanks to the period–luminosity relation which allows to measure their intrinsic brightness. They are also relatively young and bright, potentially making them excellent tracers of the young stellar population that is responsible for shaping the visible aspect of our Galaxy, the Milky Way. However, being observers embedded in the dusty interstellar medium of the Galaxy, deriving reliable photometric distances to classical Cepheids of the Milky Way is a challenge. The typical approach is to use “reddening-free” indices, such as Wesenheit magnitudes, to obviate the need for an extinction correction. However, this approach is not reliable – especially toward the inner Galaxy – as its assumption of a universal total-to-selective extinction ratio is not satisfied, particularly in lines-of-sight where the extinction is high and crosses spiral arms. We instead estimate new distances for 3425 Cepheids based on mid-IR photometry from WISE, which suffers minimally from extinction, and by adopting a 3D extinction map to calculate the necessary (albeit small) extinction corrections. We show that our distances are consistent with Gaia’s parallaxes for the subset with relative parallax errors smaller than 10%, verifying that our mean distance errors are of the order of 13% and that the mean parallax zero point for this subsample is $7\>\mu\rm{as}$. The catalog of Cepheid distances is made available online.
@@
software: TOPCAT (Taylor, 2005, 2006), Julia (Bezanson et al., 2012), Matplotlib (Hunter, 2007), IPython (Pérez & Granger, 2007), SciPy (Virtanen et al., 2020), Astropy (Astropy Collaboration et al., 2018), NumPy (Harris et al., 2020)
@@ 1 Introduction
These catalogs led to a great number of Milky Way studies, including that of the Galactic warp (Skowron et al., 2019a, b; Dehnen et al., 2023; Cabrera-Gadea et al., 2024), the spiral structure (Minniti et al., 2021; Lemasle et al., 2022; Bobylev, 2022; Gaia Collaboration et al., 2023; Semczuk et al., 2023), the rotation curve and the outer disk (Mróz et al., 2019; Ablimit et al., 2020; Drimmel et al., 2023; Bobylev & Bajkova, 2023; Beordo et al., 2024), just to name a few. All such studies require calculating distances to classical Cepheids, and in each case this is done with the use of period–luminosity (PL) relations (Leavitt, 1908). However, various studies have different approaches for the exact method of distance estimation. For example Skowron et al. (2019a, b), hereafter S19, used mid-infrared (mid-IR) data from the Wide-field Infrared Survey Explorer (WISE, Wright et al. 2010) and Spitzer Space Telescope (Werner et al., 2004), together with mid-IR PL relations and an extinction correction; Gaia Collaboration et al. (2023) used photometric data from the third Gaia data release (hereafter Gaia DR3) and period–Wesenheit–metallicity relations calibrated for the Gaia bands to avoid the extinction corrections; while Lemasle et al. (2022) use mid-IR data from WISE and period–Wesenheit relations calibrated for mid-IR WISE bands, also without an extinction correction.
@@
Figure 1 shows the comparison of distances between the study of Gaia Collaboration et al. (2023) and S19, where the distance difference is plotted against the Galactic longitude, with $l=0^{\circ}$ pointing toward the Galactic center. It is striking how large the distance differences are (despite the reported distance uncertainties of about $5-10\%$), and calls into question the detailed studies of Galactic structure with classical Cepheids. A similar comparison between distances from studies of S19 and Lemasle et al. (2022) shows discrepancies of the same order of magnitude (B. Lemasle, private communication).
@@
Interestingly, the differences in distance are very large in the direction of the Galactic center, but negligible toward the anticenter ($l=\pm 180^{\circ}$). The mean general offset of about $200$ pc between the two datasets is expected and results from the different approach to distance calculation. Instead, the distance differences in sight lines toward the Galactic center ($-100^{\circ}<l<100^{\circ}$) have a mean absolute value of about $930$ pc and a dispersion of $1400$ pc. This suggests that the main factor contributing to distance uncertainty is tied to extinction, which is highest toward the Galactic center. Indeed, the usage of the Wesenheit index in the period-Wesenheit relation to avoid the extinction correction has an underlying assumption that the ratio of total-to-selective extinction is constant throughout the Galaxy (i.e. a universal extinction curve), which is not true in regions of the disk with high levels of extinction, where the extinction curve is known to be highly variable (Fitzpatrick & Massa, 2007; Schlafly et al., 2016; Nataf et al., 2016; Maíz Apellániz & Barbá, 2018; Fouesneau et al., 2022; Zhang et al., 2023). Since a significant number of classical Cepheids are located in such regions, distance estimates based on the above assumption are less reliable toward the inner Galaxy.
@@
On the other hand, the use of the PL relation instead of the period–Wesenheit index relation requires accounting for interstellar extinction. This is typically handled by moving to the mid-IR wavelengths, where the extinction is $\sim$10 times smaller than in the optical. However, even though mid-IR extinction values are generally very low, they are non-negligible within the Galactic plane and toward the inner disk, where a large fraction of the Galaxy’s Cepheids are located. It is therefore necessary to use three-dimensional extinction maps to account for extinction. Motivated by Figure 1, we decided to derive a new set of accurate distances based on the mid-IR data, mid-IR PL relations and mid-IR extinction corrections based on 3D extinction maps, with the goal of producing a catalog of Cepheid distances that can be used for detailed Galactic structure studies.
@@ 2.1 Classical Cepheid Catalog
In summary, thanks to the combined coverage of the different surveys, we believe that our sample is highly complete to at least $G=18$ mag ($I=19.5$ mag) within the footprint of OGLE, though will be smaller in highly extincted areas toward the Galactic center. Outside this footprint our completeness will be somewhat less. Thanks to the intrinsic brightness of Cepheids and limited extinction, we are likely seeing all Milky Way Cepheids in the second and third quadrant of the Galaxy ($90^{\circ}<l<270^{\circ}$), and most Cepheids outside the Galactic bulge region (outside $30^{\circ}$ of the Galactic center) in the first and fourth quadrants, with extinction more or less limiting our view in selective directions.
Figure 3: Completeness estimate of Gaia Cepheids across galactic longitude ($l$) in 10 wide bins. Panel (a) shows the distribution of $l$ as a function of $G$ magnitude. In cyan, are all Cepheids detected with the Gaia survey, while in red, are all exclusively discovered outside of Gaia. The black dotted line indicates the magnitude limit ($84^{th}$ percentile) of Gaia Cepheids ($G<16.7$). Panel (b) shows the contribution of Cepheids from Gaia ($N_{\rm Gaia}$) to the entire catalog ($N_{all}$) as a function of $l$, restricting the sample to within the magnitude limit of Gaia, where this metric provides a measure for the completeness of Cepheids observed by Gaia. The two green dashed lines indicate completeness of 0.95 and 1.0.
@@ 2.2 Mid-IR Photometry
Moreover, for 534 bright Cepheids, we use data from the WISE All-Sky Release Source Catalog [^3] WISE All-Sky release notes in place of the AllWISE magnitudes, to avoid the problematic AllWISE measurements that were made during the Post-Cryo mission phase, as suggested by the AllWISE team444 AllWISE cautionary notes These are Cepheids brighter than $W1=$ 8.25 mag in AllWISE found in the ecliptic longitude ranges $89.4^{\circ}<\lambda<221.8^{\circ}$ and $280.6^{\circ}<\lambda<48.1^{\circ}$.
Since we wish to use the photometry from two different catalogs, it is important to confirm that the two catalogs are on the same photometric system, especially as we will use the $W1$ PL relation (Leavitt law) from Wang et al. (2018), which was calibrated using AllWISE data. Schlafly et al. (2019) recommend applying a $\Delta W1=W1_{U}-W1_{A}$ correction of 0.004 mag between the AllWISE ($W1_{A}$) and unWISE ($W1_{U}$) magnitudes, but also noted a clear magnitude dependence between the two magnitude systems (see their Figure 7). For our sample, we find an offset of $\Delta W1$ of 0.024 mag for $W1_{U}>8.25$ mag (see Figure 4), with little evidence of a magnitude term. Therefore we adopt 0.024 mag as the offset at these magnitudes to bring the unWISE photometry to the AllWISE photometric system before calculating distances.
@@
Meanwhile for magnitudes $W1_{A}<8.25$ mag we find a significantly larger $\Delta W1$ of 0.166 mag (see Figure 4). One might be tempted to interpret this as evidence that indeed the unWISE photometry should not be used at bright magnitudes, and to use only the AllWISE magnitudes. However, a more careful comparison of the photometric distances derived from AllWISE and unWISE $W1$ magnitudes revealed that, once the photometric offsets between the two catalogs are taken into account, the unWISE $W1$ distances are of similar quality to those of AllWISE, even at bright magnitudes (See section 3.1). This is perhaps to be expected since saturation will be problematic for both catalogs. Moreover, judging from the comparison to the astrometric distances based on parallax, $(\varpi+0.017~{}\mu\rm{as})^{-1}$, for the subset of 520 bright ($W1_{A}<8.25$ mag) Cepheids with good parallaxes ($\sigma_{\varpi}/\varpi<0.1$ and $ruwe<1.4$), it seems that the large $\Delta W1=0.166$ mag offset between the two catalogs at $W1<8.25$ mag is largely due to a bias in AllWISE (Figure 5). That is, the AllWISE $W1_{A}$ magnitudes at $W1_{A}<8.25$ mag are biased with respect to fainter stars in AllWISE, and it is only after adding an additional 0.142 (= 0.166 - 0.024) mag to these bright stars that the AllWISE photometric distances can be made consistent with the parallax-based distances at bright magnitudes.
@@
Figure 5: Distribution of the differences in the astrometric and photometric distances for bright ($W1_{A}<8.25$ mag) Cepheids with good parallaxes. The red histogram shows distances derived from the unWISE $W1_{U}$ magnitudes after applying an offset of 0.024 mag. The blue histograms show the same but for photometric distances derived from AllWISE $W1_{A}$ magnitudes, both with (solid line) and without (dashed line) a 0.142 offset applied to the AllWISE magnitudes.
@@
In contrast, we found that the $W2$ photometry (and distances) from AllWISE and unWISE seem to be quite consistent with each other, after adjusting for an offset of 0.024 mag, at both bright and faint magnitudes. However, the photometric distances using either set of $W2$ magnitudes, while consistent with each other, both show significant bias with respect to the astrometric distances from Gaia, which may suggest a problem in the calibration of the $W2$ PL relation. Moreover, it has been shown that the Cepheid brightness in $W2$ (4.5 $\mu{\rm m}$) is affected by carbon monoxide absorption (Marengo et al., 2010; Scowcroft et al., 2011; Monson et al., 2012) and so its use as a distance indicator has been questioned. For these reasons we have decided to proceed using only the $W1$ magnitudes.
@@
Since we wish our sample to be as complete as possible, and we found that reliable distances could be derived even for the bright problematic sources, we make no selections based on either the unWISE or AllWISE flags. However, as our sample is primarily found in the galactic plane, the photometry is potentially affected by crowding. A useful measure of potential crowding effects provided by unWISE is the quality indicator fracflux, which is the PSF-weighted fraction of flux attributed to the source at its location. A careful quantitative study of the sensitivity of fracflux to nearby sources was recently made by Saydjari et al. (2023) via experiments with injected sources, who use a similar (but improved) pipeline for the production of DECaPS2555Dark Energy Camera Plane Survey data release 2. For their images they find little or no bias in the measured photometry for fracflux $>0.5$, while the measured uncertainties are increasingly underestimated with decreasing fracflux (See lower-right panels of their Figures 21 and 22). However, since these results are based on experiments with DECaPS imaging, it would be better to explore the quality of the photometric uncertainties with respect to fracflux with unWISE data. We check for a possible contribution from crowding to the unWISE magnitude errors by cross-matching unWISE and Gaia over a large area of the sky, and comparing the unWISE $W1$ magnitudes to $W1$ magnitudes derived from Gaia photometry, whose much higher resolution will be much less affected by crowding effects. We find no significant increase in the dispersion of these magnitude differences as a function of fracflux, therefore we decided not to include any additional error terms based on this parameter. See the discussion in Appendix A for further details.

### v2

@@ Abstract
Classical Cepheids are the archetype of the standard candle, thanks to the period–luminosity relation which allows to measure their intrinsic brightness. They are also relatively young and bright, potentially making them excellent tracers of the young stellar population that is responsible for shaping the visible aspect of our Galaxy, the Milky Way. However, being observers embedded in the dusty interstellar medium of the Galaxy, deriving reliable photometric distances to classical Cepheids of the Milky Way is a challenge. The typical approach is to use “reddening-free” indices, such as Wesenheit magnitudes, to obviate the need for an extinction correction. However, this approach could lead to unknown systematics – especially toward the inner Galaxy – as its assumption of a universal total-to-selective extinction ratio is not satisfied, particularly in lines-of-sight where the extinction is high and crosses spiral arms. We instead estimate new distances for 3424 Cepheids based on mid-IR photometry from WISE, which suffers minimally from extinction, and by adopting a 3D extinction map to calculate the necessary (albeit small) extinction corrections. We show that our distances are consistent with Gaia’s parallaxes for the subset with relative parallax errors smaller than 10%, verifying that our mean distance errors are of the order of 6% and that the mean parallax zero point for this subsample is $7\>\mu\rm{as}$.
@@
software: TOPCAT (Taylor, 2005, 2006), Julia (Bezanson et al., 2012), Matplotlib (Hunter, 2007), IPython (Pérez & Granger, 2007), SciPy (Virtanen et al., 2020), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), NumPy (Harris et al., 2020)
@@ 1 Introduction
These catalogs led to a great number of Milky Way studies, including that of the Galactic warp (Skowron et al., 2019a, b; Dehnen et al., 2023; Cabrera-Gadea et al., 2024), the spiral structure (Minniti et al., 2021; Lemasle et al., 2022; Bobylev, 2022; Gaia Collaboration et al., 2023; Semczuk et al., 2023), the rotation curve and the outer disk (Mróz et al., 2019; Ablimit et al., 2020; Drimmel et al., 2023; Bobylev & Bajkova, 2023; Beordo et al., 2024), just to name a few. All such studies require calculating distances to classical Cepheids, and in each case this is done with the use of period–luminosity (PL) relations (Leavitt, 1908). However, various studies have different approaches to the exact method of distance estimation. For example Skowron et al. (2019a, b), hereafter S19, used mid-infrared (mid-IR) data from the Wide-field Infrared Survey Explorer (WISE, Wright et al. 2010) and Spitzer Space Telescope (Werner et al., 2004), together with mid-IR PL relations and an extinction correction; Gaia Collaboration et al. (2023) used photometric data from the third Gaia data release (hereafter Gaia DR3) and period–Wesenheit–metallicity relations calibrated for the Gaia bands to avoid the extinction corrections; while Lemasle et al. (2022) use mid-IR data from WISE and period–Wesenheit relations calibrated for mid-IR WISE bands, also without an extinction correction.
@@
Figure 1 shows the comparison of distances between the study of Gaia Collaboration et al. (2023) and S19, where the distance difference is plotted against the Galactic longitude, with $l=0^{\circ}$ pointing toward the Galactic center. It is striking how large the distance differences are (despite the reported distance uncertainties of about $5-10\%$), and calls into question the detailed studies of Galactic structure with classical Cepheids. A similar comparison between distances from studies of S19 and Lemasle et al. (2022) shows discrepancies of the same order of magnitude, though not necessarily the same sign (B. Lemasle, private communication).
@@
Interestingly, the differences in distance (Figure 1) are very large in the direction of the Galactic center, but negligible toward the anticenter ($l=\pm 180^{\circ}$). The mean general offset of about $200$ pc between the two datasets is expected and results from the different approach to distance calculation. Instead, the distance differences in sight lines toward the Galactic center ($-100^{\circ}<l<100^{\circ}$) have a mean absolute value of about $930$ pc and a dispersion of $1400$ pc. This suggests that the main factor contributing to distance uncertainty is tied to extinction, which is highest toward the Galactic center. Indeed, the usage of the Wesenheit index in the period-Wesenheit relation to avoid the extinction correction has an underlying assumption that the ratio of total-to-selective extinction is constant throughout the Galaxy (i.e. a universal extinction curve), which is not true in regions of the disk with high levels of extinction, where the extinction curve is known to be highly variable (Fitzpatrick & Massa, 2007; Schlafly et al., 2016; Nataf et al., 2016; Maíz Apellániz & Barbá, 2018; Fouesneau et al., 2022; Zhang et al., 2023; Zhang & Green, 2025). Since a significant number of classical Cepheids are located in such regions, it is preferable to use distance estimates less sensitive to the assumption of a universal extinction curve.
@@
On the other hand, the use of the PL relation instead of the period–Wesenheit index relation requires accounting for interstellar extinction. This is typically handled by moving to the mid-IR wavelengths, where the extinction is $\sim$10 times smaller than in the optical. However, even though mid-IR extinction values are generally very low, they are non-negligible within the Galactic plane and toward the inner disk, where a large fraction of the Galaxy’s Cepheids are located. It is therefore necessary to use three-dimensional extinction maps to account for extinction. Motivated by Figure 1, we decided to derive a new set of accurate distances based on the mid-IR data, mid-IR PL relations and mid-IR extinction corrections from 3D extinction maps, with the goal of producing a catalog of Cepheid distances that can be used for detailed Galactic structure studies.
@@ 2.1 Classical Cepheid Catalog
Figure 3: Completeness estimate of Gaia  Cepheids across Galactic longitude ($l$) in 10 wide bins. Panel (a) shows the distribution of $l$ as a function of $G$ magnitude. In cyan, are all Cepheids detected with the Gaia survey, while in red, are all exclusively discovered outside of Gaia. The black dotted line indicates the magnitude limit ($84^{th}$ percentile) of Gaia Cepheids ($G<16.7$ mag). Panel (b) shows the contribution of Cepheids from Gaia ($N_{\rm Gaia}$) to the entire catalog ($N_{all}$) as a function of $l$, restricting the sample to within the magnitude limit of Gaia, where this metric provides a measure for the completeness of Cepheids observed by Gaia. The two green dashed lines indicate completeness of 0.95 and 1.0.
@@
In summary, thanks to the combined coverage of the different surveys, we believe that our sample is highly complete to at least $G=18$ mag ($I=19.5$ mag) within the footprint of OGLE, though will be smaller in highly extincted areas toward the Galactic center. Outside this footprint our completeness will be somewhat lower. Thanks to the intrinsic brightness of Cepheids and limited extinction, we are likely seeing all Milky Way Cepheids in the second and third quadrant of the Galaxy ($90^{\circ}<l<270^{\circ}$), and most Cepheids outside the Galactic bulge region (outside $30^{\circ}$ of the Galactic center) in the first and fourth quadrants, with extinction more or less limiting our view in selective directions.
@@ 2.2 Mid-IR Photometry
Moreover, for 534 bright Cepheids, we use data from the WISE All-Sky Release Source Catalog [^3] WISE All-Sky release notes in place of the AllWISE magnitudes, to avoid the problematic AllWISE measurements that were made during the Post-Cryo mission phase, as suggested by the AllWISE team444 AllWISE cautionary notes. These are Cepheids brighter than $W1=$ 8.25 mag in AllWISE found in the ecliptic longitude ranges $89.4^{\circ}<\lambda<221.8^{\circ}$ and $280.6^{\circ}<\lambda<48.1^{\circ}$.
Since we wish to use the photometry from two different catalogs (unWISE and AllWISE), it is important to confirm that the two catalogs are on the same photometric system, especially as we will use the $W1$ PL relation (Leavitt law) from Wang et al. (2018), which was calibrated using AllWISE data. Schlafly et al. (2019) recommend applying a $\Delta W1=W1_{U}-W1_{A}$ correction of 0.004 mag between the AllWISE ($W1_{A}$) and unWISE ($W1_{U}$) magnitudes, but also noted a clear magnitude dependence between the two magnitude systems (see their Figure 7). For our sample, we find an offset $\Delta W1$ of 0.024 mag for $W1_{U}>8.25$ mag (see Figure 4), with little evidence of a magnitude term. Therefore we adopt 0.024 mag as the offset at these magnitudes to bring the unWISE photometry to the AllWISE photometric system before calculating distances.
@@
Meanwhile for magnitudes $W1_{A}<8.25$ mag we find a significantly larger $\Delta W1$ of 0.166 mag (see Figure 4). One might be tempted to interpret this as evidence that indeed the unWISE photometry should not be used at bright magnitudes, and to use only the AllWISE magnitudes. However, a more careful comparison of the photometric distances derived from AllWISE and unWISE $W1$ magnitudes revealed that, once the photometric offsets between the two catalogs are taken into account, the unWISE $W1$ distances are of similar quality to those of AllWISE, even at bright magnitudes (See section 3.1). This is perhaps to be expected since saturation will be problematic for both catalogs. Moreover, judging from the comparison to the astrometric distances based on parallax, $(\varpi+0.017~{}\mu\rm{as})^{-1}$, for the subset of 520 bright ($W1_{A}<8.25$ mag) Cepheids with good parallaxes ($\sigma_{\varpi}/\varpi<0.1$ and $ruwe<1.4$), it seems that the large $\Delta W1=0.166$ mag offset between the two catalogs at $W1<8.25$ mag is largely due to a bias in AllWISE (see upper panel in Figure 5). That is, the AllWISE $W1_{A}$ magnitudes at $W1_{A}<8.25$ mag are biased with respect to fainter stars in AllWISE, and it is only after adding an additional $0.142$ $(=0.166-0.024)$ mag to these bright stars that the AllWISE photometric distances can be made consistent with the parallax-based distances at bright magnitudes.
@@
Figure 5: Distribution of the differences in the astrometric and photometric distances for bright ($W1_{A}<8.25$ mag) Cepheids with good parallaxes using $W1$ (top panel) and $W2$ (bottom panel). The red histogram shows differences derived from the unWISE magnitudes and the blue from the AllWISE magnitudes. An offset of 0.024 mag has been applied to the unWISE magnitudes, while the differences between the astrometric and photometric distances derived from AllWISE $W1_{A}$ magnitudes are shown both with (solid line) and without (dashed line) a 0.142 offset applied to the AllWISE magnitudes for stars brighter than $W1<8.25$.
@@
In contrast, we found that the $W2$ photometry (and distances) from AllWISE and unWISE seem to be quite consistent with each other, after adjusting for an offset of 0.024 mag, at both bright and faint magnitudes. However, the photometric distances using either set of $W2$ magnitudes, while consistent with each other, both show significant bias with respect to the astrometric distances from Gaia (see lower panel in Figure 5). This bias is likely due to an unknown offset in the $W2$ magnitudes present in both catalogs at bright magnitudes. Moreover, it has been shown that the Cepheid brightness in $W2$ (4.5 $\mu{\rm m}$) is affected by carbon monoxide absorption (Marengo et al., 2010; Scowcroft et al., 2011; Monson et al., 2012) and so its use as a distance indicator has been questioned. For these reasons we have decided to proceed using only the $W1$ magnitudes.
2.2.1 Crowding
@@
Since we wish our sample to be as complete as possible, and we found that reliable distances could be derived even for the bright problematic sources, we make no selections based on either the unWISE or AllWISE flags. However, as our sample is primarily found in the Galactic plane, and given that the PSF of WISE $W1$ is large (6.1”) [^5] https://wise2.ipac.caltech.edu/docs/release/allsky/, it is possible that in the most crowded regions of the Galaxy the WISE photometry is potentially affected by crowding. A useful measure of potential crowding effects provided by unWISE is the quality indicator fracflux, which is the PSF-weighted fraction of flux attributed to the source at its location. A careful quantitative study of the sensitivity of fracflux to nearby sources was recently made by Saydjari et al. (2023) via experiments with injected sources, who use a similar (but improved) pipeline for the production of DECaPS2666Dark Energy Camera Plane Survey data release 2. For their images they find little or no bias in the measured photometry for fracflux $>0.5$, while the measured uncertainties are increasingly underestimated with decreasing fracflux (See lower-right panels of their Figures 21 and 22). However, since these results are based on experiments with DECaPS imaging, it would be better to explore the quality of the photometric uncertainties with respect to fracflux with unWISE data. We check for a possible contribution from crowding to the unWISE magnitude errors by cross-matching unWISE and Gaia over a large area of the sky, and comparing the unWISE $W1$ magnitudes to $W1$ magnitudes derived from Gaia photometry, whose much higher resolution will be much less affected by crowding effects. We find no significant increase in the dispersion of these magnitude differences as a function of fracflux. See the discussion in Appendix A for further details.
However, the above analysis was necessarily conducted over an area of the sky with little or no extinction, outside the Galactic plane, where crowding is minimal. Therefore it only tests the effect on the photometry when there is a nearby, possibly blended source, but does not test for the additional biases that may come from a noisier background present in a crowded field. In order to check this, we compare our $W1$ photometry with Spitzer [3.6] photometry as these passbands are quite similar, while Spitzer has a much smaller PSF of 1.66” (Fazio et al., 2004). Specifically, we cross-match within a one arcsecond search radius our Cepheids with the Galactic Legacy Infrared Midplane Survey Extraordinaire (GLIMPSE) catalogue, including the extensions of their survey: GLIMPSE II, GLIMPSE 3D, GLIMPSE 360, Deep GLIMPSE, Vela-Carina, GLIMPSE Proper, APOGLIMPSE, ”Spitzer Mapping of the Outer Galaxy” (SMOG) and ”A Spitzer Legacy Survey of the Cygnus-X Complex” (Cygnus-X) programs (Benjamin et al., 2003; Churchwell et al., 2009). We recover [3.6] magnitudes for 2019 Cepheids within the Glimpse footprint. The $W1$ and [3.6] passbands are quite similar, so we perform a straight difference, ($W1-$[3.6]). We find that the median difference is 0.012 magnitudes (RMS = 0.17), which decreases to $-0.0015$ magnitudes (RMS = 0.13) if we restrict to sources fainter than [3.6] $>$ 8 mag. In general we do see higher ($W1-$[3.6]) differences (i.e. scatter) toward the inner Galaxy (RMS = 0.25 for Cepheids within $50^{\circ}$ of the Galactic center), where crowding can be an issue. To directly test the effect of crowding on the photometry, we tile the Glimpse compilation using the HEALPix pixelation scheme (Górski et al., 2005) at level 6 (Nside = 64), resulting in an area per pixel of 0.84 square degrees. We counted the number of Glimpse sources within each HEALPix to derive the source density per square degree at the $W1$/[3.6] wavelengths. The ($W1-$[3.6]) difference with respect to the number density in each HEALPix, shown in Figure 6, confirms that there is a slight increase in the RMS with respect to the source density of the field, but there is no significant bias. It is also worth noting, that some of the above scatter is a result of the stellar variability. The mean brightness of a Cepheid is based on about 10 epochs in the case of WISE, while typically on 2 epochs in the case of Spitzer. This means that the uncertainty of estimating a mean Cepheid magnitude from Spitzer is over twice the uncertainty from WISE (for details see Section 2.2.2). We therefore conclude that crowding has no substantial effect on the $W1$ magnitudes, and for simplicity decided not to include any additional error terms.
Figure 6: The difference between the WISE $W1$ (averaged) and Spitzer [3.6] $\mu{\rm m}$ magnitudes for the Cepheids in this study, shown as a function of source density per square degree (blue points). The running median and the 16th and 84th percentiles are shown as red solid and dashed lines, respectively.
2.2.2 Photometric uncertainties